A department store manager has monitored the number of complaints received per week about poor service. The probabilities for numbers of complaints in a week, established by this review, are shown in the table. Number of complaints 0 1 2 3 4 5 Probability 0.20 0.29 0.26 0.12 0.05 0.08 What is the probability of between 1 and 3 (inclusive) complaints received per week? Please specify your answer in decimal terms and round your answer to the nearest hundredth (e.g., enter 12 percent as 0.12). Note that the correct answer will be evaluated based on the full-precision result you would obtain using Excel.
Answer #1

Let the random variable X denotes the number of complaints in week Given that the probability distribution of X is X=u P(x=2)probabil Required ility pliex=3) P(x=1) + P(x=2) 2 + P(x=3) 0.26 + 0.12 0.29 0.67 =